OpenStudy (anonymous):

Factor completely: ax + 2ay + bx + 2by @ash2326

OpenStudy (ash2326):

We have \[ax+2ay+bx+2by\] Let's take out a from the first two terms, we get \[a(x+2y)+bx+2by\] @unheard can you take out anything common from the last two terms?

OpenStudy (anonymous):

b is common.

OpenStudy (ash2326):

Can you take it out?

OpenStudy (anonymous):

I forgot how .-.

OpenStudy (ash2326):

\[ b\times x +2\times b \times y\] b is multiplied to both the terms. Let's divide both terms by b we get \[(x+2y)\] do you understand it ?

OpenStudy (anonymous):

Just.. take em out? .-. Sorry. I'm not good at this at all. > . <

OpenStudy (ash2326):

no problem, we'll learn it.

OpenStudy (ash2326):

ok, for example we have \[x+2xy\] can you try?

OpenStudy (anonymous):

divide both sides by x?

OpenStudy (ash2326):

yes, divide by x and bring it out

OpenStudy (anonymous):

2y?

OpenStudy (ash2326):

Ok, I'll show how to divide \[\frac{x}{x}+\frac{2xy}{x}\] \[\frac{1 \cancel x}{\cancel x}+\frac{2y \cancel x}{\cancel x}\] We get 1+2y

OpenStudy (anonymous):

Oh, I thought it'd be 1 but I wasn't sure.

OpenStudy (ash2326):

no problem. So we get \[x(1+2y)\] Do you get it?

OpenStudy (anonymous):

Yes, I get it now.

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@radar

OpenStudy (ash2326):

sorry i was away

OpenStudy (ash2326):

Are you here?

OpenStudy (anonymous):

@radar helped me with the rest. But thanks!

OpenStudy (ash2326):

Cool, let me know if you need help with anything else.